# Fibred surfaces and their unitary rank

### v1

@@ Abstract
Let $f\colon S\to B$ a fibred surface with fibres of genus $g\geq 2$ and base curve of genus $b$. Let $u_{f}$ be its unitary rank. We prove many new slope inequalities involving $u_{f}$ and some other invariants of the fibration. As applications:
(1) we prove a new Xiao-type bound on $u_{f}$ with respect to $g$ for non-isotrivial fibrations:
@@
In particular this imples that if $f$ is not locally trivial and $u_{f}=g-1$ is maximal, then $g\leq 6$;
(2) we prove a result in the direction of the Coleman-Oort conjecture: a new constrain on the rank of the (-1,0) part of the maximal unitary Higgs subbundle of a curve generically contained in the Torelli locus;
(3) we study in the detail the extremal case (for non-isotrivial $f$) where $u_{f}=g-1$, giving many constrains for the case $g=6$, $u_{f}=5$.
2010 Mathematics Subject Classification. 14J10, 14D06, 14D07, 14G35, 14H40. Key words and phrases: fibred varieties, slope inequality, families of Jacobians, Hodge bundle, Coleman-Oort conjecture
@@ 1 Introduction
This paper deals with the geography of fibred surfaces and some applications. The main results can be divided in four sections, which we list in order of appearance.
New slope inequalities
Let $f\colon S\to B$ a relatively minimal fibred surface. We prove new slope inequalities, i.e. bounds of the form
@@
where $\alpha$ is a function of some invariants of the fibration. In particular, in our $\alpha$ will be a function depending on the genus $g$ of the general fibre, on the unitary rank $u_{f}$, and in some cases on the Clifford index of the general fibre $c_{f}$, and on the rank $r_{1}$ of the first subbundle of the Harder-Narasimhan filtration of the Hodge bundle $f_{*}\omega_{f}$ (see Section 2 for the definitions). Some of the most significant inequalities we prove are the following:
@@
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$.
@@
    Moreover, equality holds in (1.2) if and only if $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$, where $\Gamma$ is a section of $f$.
@@
    Moreover, if equality holds in (1.3), then
    • if $u_{f}\leq g-2$ then $F$ is hyperelliptic;
    • if $u_{f}=g-1$ then $F$ is trigonal.
@@
A new Xiao type inequality for the unitary rank
Using inequality (4.3) we extend to the unitary rank the bound obtained by Konno in [34] that was itself a sharpening of the famous Xiao’s bound [52]. In particular, we obtain that:
@@
    if the unitary rank $u_{f}=g-2$, then necessarily $g\leq 11$.
This is, up to now, the sharpest known bound for $u_{f}$; see Section 2 for the known results and conjectures.
Results on the case with high unitary rank
Recall that Pirola proved in [47] that for a relatively minimal fibration such that $u_{f}=g-1$, then up to a base change $q_{f}=u_{f}$. Also the following result is essentially due to Pirola (in a private communication with the author) and to the already known results [6], [30].
Theorem 1.8 (Proposito 5.1, Theorem 5.2).
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then $f$ is a non-isotrivial trigonal fibration, the smooth fibres $F_{t}$ are a covering of a non-isotrivial family of elliptic curves. Moreover, $f$ is not not Kodaira, i.e. $K_{f}^{2}<12\chi_{f}$.
Applications to the Coleman-Oort conjecture
Having proved these sharper inequalities, we can apply them to the techniques developed by Viehweg and Zuo and then by Chen, Lu and Zuo (see [50], [17], [40]) and give a result that goes in the direction of the Coleman-Oort conjecture on the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of genus $g$ (see Section 7 for the definitions and some explanation):
@@
This is a sharpening of a theorem of Chen, Lu and Zuo [17, Thm 1.1.2]: see Remark 7.5.
Results on the case with $g=6$, $u_{f}=5$
Eventually we can use our arguments, together with some other result on trigonal curves, to give some information on the extremal cases. First recall that Beorchia, Pirola and Zucconi proved in [13] that if $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $6$ and with $q_{f}=5$ then $F$ is trigonal of special Maroni invariant.
Theorem 1.10 (Proposition 8.1, Corollary 8.8).
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $6$ and with $u_{f}=5$. Then:
  1. 1.
    the surface $S$ has positive index $\tau(S)>0$.
  2. 2.
    The schematic base locus associated to the morphism $f^{*}\mathcal{A}\to\omega_{f}$ has a horizontal component $G$ of multiplicity at least 5.
  3. 3.
    If the multiplicity of $G$ is at least $6$ then $G$ is a section and the point $G\cap F$ on the general fibre is a Weierstrass point and we can list its possible gap sequences.
  4. 4.
    We can track down the possibilities for the numerical class of $K_{f}$, reducing them to the following cases:
    • (a)
      $K_{f}\equiv 10\Gamma+\mu_{1}F$;
    • (b)
      $K_{f}\equiv 9\Gamma+\Gamma^{\prime}+\mu_{1}F$;
    • (c)
      $K_{f}\equiv 6\Gamma+4\Gamma^{\prime}+\mu_{1}F$;
    • (d)
      • either $K_{f}\equiv 5\Gamma+5\Gamma^{\prime}+\mu_{1}F$,
      • or $K_{f}\equiv 5\Sigma$
    where $\Gamma$ and $\Gamma^{\prime}$ are distinct sections of $f$ and $\Sigma$ is a bisection of $f$.
Remark 1.11.
So we have given many constrains to the case of non locally trivial fibrations with $g=6$, $u_{f}=5$, which add up to the ones already given by Beorchia, Pirola, Zucconi [13], and also to the explicit conditions found by Colombo, Frediani and Pirola in the recent paper [19, Section 10]. Still, the existence of such fibred surfaces is not known, the conjecture being that they do not exist. On the other hand, for $g\leq 4$ these fibration exists -the first example being given by Pirola in [46]- and have been thoroughly investigated ([1], [44], [18]).
Moreover, we can find some conditions for the case $g=5$ and $u_{f}=4$, listed in Proposition 8.17.
For these last results we develop a finer slope inequality, applicable to these extremal cases (Proposition 6.4).
A few words about the proofs. We work in the framework of the celebrated Xiao’s method; however, a new key ingredient is a re-elaboration of a result due to Konno in [34], where he gives an original construction of a nef divisor on the surface. Moreover, we use in many points the algebraic index theorem, in particular to give characterizations of the extremal cases. In the last section, with an ad-hoc argument, we prove a very strong slope inequality holding in the extremal cases $u_{f}=g-1$. This, together with the use of results of Coppens ([20, 21, 22]) and Brundu-Sacchiero ([14]) allows us to give the results about this extremal case.
Acknowledgements This article would not have been written without the encouragement and stimulation of Gian Pietro Pirola, whom I thank heartily.
@@ 2 Fibred surfaces and their geography
We call fibred surface or sometimes simply fibration the data of a morphism $f\colon S\rightarrow B$ from a smooth projective surface $S$ to a smooth projective curve $B$ which is surjective with connected fibres. We denote with $b=g(B)$ the genus of the base curve. A general fibre $F$ is a smooth curve and its genus $g=g(F)$ is by definition the genus of the fibration. From now on, we consider fibrations of genus $g\geq 2$.
We say that $f$ is relatively minimal if it does not contain any $(-1)$-curves in its fibres. This condition is equivalent to $K_{f}$ being a relatively nef divisor. It indeed holds that $K_{f}$ is nef ,as proved for instance in [5]. Given any fibration of genus $g\geq 2$ we can contract all the $(-1)$-curves contained in the fibres and obtain a unique relatively minimal genus $g$ fibration $\overline{f}\colon\overline{S}\to B$. We will always assume, unless differently stated, that a fibred surface is relatively minimal.
@@
  • $e_{f}:=e(S)-e(B)e(F)=e(S)-4(g-1)(b-1)$ the relative topological characteristic (with $e(X)$ topological characteristic of $X$);
@@
  • We can also consider as invariants of the fibration the gonality and the Clifford index of $F$ which we denote $gon(F)$ and $c_{f}=\mbox{{Cliff}}(F)$ respectively.
From Groethendieck-Riemann-Roch theorem we have Noether’s relation
@@
We say that a fibred surface is: smooth if every fibre is smooth; isotrivial if all smooth fibres are mutually isomorphic; locally trivial if $f$ is smooth and isotrivial (equivalently $f$ is a fibre bundle by the Grauert-Fisher Theorem [9, Theorem I.10.1]); trivial if $S$ is birationally equivalent to $F\times B$ and $f$ corresponds to the projection on $B$. If $b>0$ and $f$ is relatively minimal this is equivalent to $S=F\times B$. Eventually, we say that $f$ is a Kodaira fibration if it is smooth and non isotrivial.
Let us first summarize what we know about the non-negativity of these invariants.
@@
Let $f\colon S\to B$ a relatively minimal fibred surface. The following results hold:
@@
    $K_{f}^{2}\geq 0$ and $K_{f}^{2}=0$ if and only if $f$ is locally trivial (see [5] and Remark 2.16);
@@
    $\chi_{f}\geq 0$ and $\chi_{f}=0$ if and only if $f$ is locally trivial (see [9]);
@@
    $q_{f}\leq g$ and equality holds if and only if $f$ is trivial (see [12] and also Remark 2.9).
@@
By using Grothendieck-Riemann-Roch theorem or Leray’s spectral sequence, we see that $\deg f_{*}\omega_{f}=\chi_{f}$ for any fibration $f$.
@@
A vector bundle $\mathcal{E}$ over a smooth curve is called nef if one of the following equivalent properties holds: (1) the corresponding tautological sheaf $\mathcal{O}_{\mathbb{P}(\mathcal{E})}(1)$ is nef; (2) if all its quotients have non-negative degree.
@@
The Hodge bundle of a fibration $f\colon S\to B$ of genus $g\geq 1$ is a rank $g$ nef vector bundle over $B$ ([33]).
The Hodge bundle can be decomposed in two ways as a direct summand of vector sub-bundles as follows:
@@
    where $\mathcal{A}$ ample and $\mathcal{U}$ unitary flat.
@@
So, we see that $q_{f}$ is the rank of the biggest trivial subbundle of $\mathcal{E}$, moreover, we see that $\chi_{f}=\deg\mathcal{G}=\deg\mathcal{A}$. We can also see directly from the first decomposition that $q_{f}\leq g$ and that if $q_{f}=g$ then $\chi_{f}=0$, and so $f$ is locally trivial (but Beauville’s result (4) of Theorem 2.2 is stronger).
@@
and then it holds that $q_{f}\leq u_{f}$. Observe that $\deg\mathcal{U}=0$, so we have that $\chi_{f}=\deg\mathcal{A}$, and moreover, $\mathcal{U}$ is the biggest degree 00 sub-bundle of $f_{*}\omega_{f}$, so $u_{f}=g$ if and only if $\chi_{f}=0$ (equivalently $f$ is locally trivial): compare with (2) of Theorem 2.2.
One of the most important areas of study in this field is the so called geography of fibred surfaces, i.e. the study of the inequalities holding between the invariants. We will list three kinds of these inequalities. For very nice expositions of these kind of problems see also [42] and [10].
Inequalities between the invariants: bounds on the ratio $K_{f}^{2}/\chi_{f}$
First of all, from Noether’s relation (2.1) and ((3)) of Theorem 2.2 we have:
@@
Moreover, if $f$ is relatively minimal and equality holds, then $f$ is a Kodaira fibration.
@@
Given any relatively minimal non-locally trivial fibred surface $f\colon S\to B$ over a base of genus $b\geq 2$, the surface $S$ is minimal and of general type. So, Bogomolov-Miyaoka-Yau inequality holds: $K_{S}^{2}\leq 9\chi(\mathcal{O}_{S})$. This on the relative invariants implies that
$K_{f}^{2}\leq 9\chi_{f}+(g-1)(b-1),$
which is not a “slope inequality” in the sense made explicit in the introduction, nevertheless it gives interesting informations.
@@
Note moreover that the ratio 8 is particularly significant for the geometry of the surface. Indeed, consider the topological index $\tau(S)$ of the surface, i.e. the index of the intersection form on $H^{2}(S,\mathbb{C})$ (the difference between the number of its positive and negative eigenvalues). We have that $\tau(S)=\frac{1}{3}(K_{S}^{2}-2c_{2}(S))$. So,
$K_{f}^{2}-8\chi_{f}=K_{S}^{2}-8\chi(\mathcal{O}_{S})=K_{S}^{2}-2c_{2}(S)=3\tau(S).$
For what concerns lower bounds on the ratio $K_{f}^{2}/\chi_{f}$, the first and most famous result is the so-called canonical slope inequality:
@@
[Xiao [51], Cornalba-Harris [23, Sto], see also [4, Thm. 4.5], Konno [35]]
@@
If $f$ is non-locally trivial and the equality in (2.15) is reached, then $F$ is hyperelliptic and $q_{f}=0$.
@@
Suppose that $K_{f}^{2}=0$. Then by the slope inequality we have $\chi_{f}=0$ so $f$ is locally trivial. If, on the other hand, $f$ is locally trivial; then by $(ii)$ we have $\chi_{f}=0$, then by Noether’s relation and the non-negativity of $e_{f}$ we have $K_{f}^{2}=0$.
Remark 2.17.
Even from the classical results of Theorem 2.14, it is natural to ask for bound increasing with $c_{f}$ and with $q_{f}$. Many authors have contributed to this topic, here we just recall [36], [7], [8], [37], [39].
Inequalities between the invariants: Xiao’s type of bounds
As mentioned in $q_{f}\leq g$, and equality holds if and only if $f$ is trivial. So, it is natural to ask: if $f$ is not trivial, what is the inequality between $g$ and $q_{f}$? From now on we assume $f$ is non-trivial. Xiao initiated the study in [52] by proving that if $b=0$

### v2

@@ Abstract
Let $f\colon S\to B$ a complex fibred surface with fibres of genus $g\geq 2$. Let $u_{f}$ be its unitary rank, i.e., the rank of the maximal unitary summand of the Hodge bundle $f_{*}\omega_{f}$. We prove many new slope inequalities involving $u_{f}$ and some other invariants of the fibration. As applications:
  • (1)
    we prove a new Xiao-type bound on $u_{f}$ with respect to $g$ for non-isotrivial fibrations:
@@
    In particular this implies that if $f$ is not locally trivial and $u_{f}=g-1$ is maximal, then $g\leq 6$;
  • (2)
    we prove a result in the direction of the Coleman-Oort conjecture: a new constraint on the rank of the $(-1,0)$ part of the maximal unitary Higgs subbundle of a curve generically contained in the Torelli locus.
2010 Mathematics Subject Classification. 14J10, 14D06, 14D07, 14G35, 14H40.
Key words and phrases: fibred varieties, slope inequality, families of Jacobians, Hodge bundle, Coleman-Oort conjecture.
@@ 1 Introduction
This paper deals with the geography of fibred surfaces and some applications. We work over the complex field $\mathbb{C}$. The main results can be divided in four sections, which we list in order of appearance.
1.1 New slope inequalities
Let $f\colon S\to B$ be a relatively minimal fibred surface. The relative canonical line bundle is $\omega_{f}:=\omega_{S}\otimes f^{*}\omega_{B}^{-1}$; we call $K_{f}$ any associated divisor. Let $F$ be a general fibre of $f$. Let $\chi_{f}:=\chi(\mathcal{O}_{S})-\chi(\mathcal{O}_{F})\chi(\mathcal{O}_{B})$ be the relative Euler characteristic (see Section 2).
We prove new slope inequalities, i.e., bounds of the form
@@ 1.1 New slope inequalities
where $\alpha$ is a function of some invariants of the fibration. In particular, in our case $\alpha$ will be a function depending on the genus $g$ of the general fibre, on the unitary rank $u_{f}$, and in some cases on the Clifford index of the general fibre $c_{f}$. The most significant inequalities we prove are the following:
@@
Let $f\colon S\to B$ be a relatively minimal locally non-trivial fibred surface of genus $g\geq 2$.
@@
    Moreover, the equality holds in (1.2) if and only if $u_{f}=g-1$, the general fibre $F$ is trigonal and there exists a section $\Gamma$ of $f$ such that $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$.
@@
    Moreover, if the equality holds in (1.3), then necessarily $u_{f}=g-1$ and $F$ is trigonal.
@@
1.2 A new Xiao type inequality for the unitary rank
We extend the bound obtained by Konno in [21] to the unitary rank. Konno’s bound was itself a sharpening of the famous Xiao’s bound [42]. Moreover, our bound is strict. In particular, we obtain that:
@@ 1.2 A new Xiao type inequality for the unitary rank
    if the unitary rank is $u_{f}=g-2$, then necessarily $g\leq 11$.
The bound (1.7) is, to our knowledge, the sharpest known bound for $u_{f}$; see Section 2 for the known results and conjectures.
1.3 Results on the case with high unitary rank
Pirola proved in [35] that, for a fibration such that $u_{f}=g-1$, up to an étale base change the relative irregularity coincides with the unitary rank $q_{f}=u_{f}$. Also the following result is essentially due to Pirola (in a private communication with the author) and to the already known results [3], [19].
Theorem 1.8 (Proposition 5.1, Theorem 5.2).
Let $f\colon S\to B$ a relatively minimal locally non-trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then $f$ is a non-isotrivial trigonal fibration and its smooth fibres $F_{t}$ are a covering of a non-isotrivial family of elliptic curves $E_{t}$. Moreover, $f$ is not Kodaira, i.e., $K_{f}^{2}<12\chi_{f}$.
1.4 Applications to the Coleman-Oort conjecture
We apply the inequalities of Theorem 1.1 to the study of Shimura varieties on the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of genus $g$. We work in the framework established by Viehweg and Zuo in [40] and developed by Chen, Lu and Zuo in [10] and in [26]. We obtain a result that goes in the direction of the Coleman-Oort conjecture. See Section 7 for the notation and some explanation.
@@ 1.4 Applications to the Coleman-Oort conjecture
This is a sharpening of a theorem of Chen, Lu and Zuo [10, Thm 1.1.2]: see Remark 7.5.
A few words about the proofs. We work in the framework of the celebrated Xiao’s method, with at least two new key ingredients: first we improve a result due to Konno in [21] that gives an original construction of a nef divisor on the surface (see also Remark 3.13). Here we push Konno’s idea to his full capacity, connecting the results with the unitary rank of the fibration, while in his paper he only dealt with the relative irregularity. Our approach is more natural, as the unitary flat subbundle in the second Fujita decomposition sits naturally in the Harder-Narasimhan filtration of the Hodge bundle: see Remark 3.4. Moreover, to obtain the strong inequalities results in Theorem 1.1, we use crucially the new Clifford-type inequality for canonical curves proved by the author together with Riva in [36].
With Xiao’s approach it is very hard to prove something on the extremal cases: for example the fact that surfaces attaining the classical slope (2.16) are hyperelliptic cannot be derived from the Xiao method. In Theorem 1.1, instead, we give much information on the cases reaching the equalities. This is done using other techniques: in particular some results in [19], the algebraic index theorem and results of Lu and Zuo [27] and Konno [22].
Acknowledgements This article would not have been written without the encouragement and stimulation of Gian Pietro Pirola, whom I thank heartily. I thank the referees for their careful reading of the paper and for many useful suggestions, in particular I owe to one of them a sharpening of Theorem 1.1. I am partially supported by the PRIN project 20228JRCYB “Moduli spaces and special varieties” and by GNSAGA - INdAM.
@@ 2 Fibred surfaces and their geography
We call a fibred surface or sometimes simply a fibration the data of a morphism $f\colon S\rightarrow B$ from a smooth projective surface $S$ to a smooth projective curve $B$ which is proper, surjective and with connected fibres. We denote by $b=g(B)$ the genus of the base curve. A general fibre $F$ is a smooth curve and its genus $g=g(F)$ is by definition the genus of the fibration. From now on, we consider fibrations of genus $g\geq 2$.
We say that $f$ is relatively minimal if it does not contain any $(-1)$-curves in its fibres. This condition is equivalent to the relative canonical divisor $K_{f}=K_{S}-f^{*}K_{B}$ being relatively nef. Given any fibration of genus $g\geq 2$ we can contract all the $(-1)$-curves contained in the fibres and obtain a unique relatively minimal genus $g$ fibration $\overline{f}\colon\overline{S}\to B$.
@@
  • $e_{f}:=e(S)-e(B)e(F)=e(S)-4(g-1)(b-1)$ the relative topological characteristic (with $e(S)=c_{2}(S)$ the topological characteristic of $S$);
@@
  • the gonality $gon(F)$ of the general fibre and its Clifford index $c_{f}=\mbox{{Cliff}}(F)$.
From the Groethendieck-Riemann-Roch theorem we have the relative Noether equality:
@@
We say that a fibred surface is smooth if every fibre is smooth; isotrivial if all smooth fibres are mutually isomorphic; locally trivial if $f$ is smooth and isotrivial (equivalently $f$ is a fibre bundle by the Grauert-Fisher Theorem); trivial if $S$ is birationally equivalent to $F\times B$ and $f$ corresponds to the projection on $B$. If $f$ is relatively minimal this is equivalent to $S=F\times B$. Eventually, we say that $f$ is a Kodaira fibration if it is smooth and non-isotrivial.
Let us summarize what we know about the non-negativity of these invariants.
@@
Let $f\colon S\to B$ be a fibred surface. The following results hold:
@@
    Suppose that $f$ is relatively minimal. Then $K_{f}^{2}\geq 0$ and $K_{f}^{2}=0$ if and only if $f$ is locally trivial;
@@
    $\chi_{f}\geq 0$ and $\chi_{f}=0$ if and only if $f$ is locally trivial;
@@
    $q_{f}\leq g$ and the equality holds if and only if $f$ is trivial (see [7] and also Remark 2.9).
@@
By using the Grothendieck-Riemann-Roch theorem or Leray’s spectral sequence, we see that $\deg f_{*}\omega_{f}=\chi_{f}$ for any fibration $f$.
@@
A vector bundle $\mathcal{E}$ over a smooth curve is called nef if one of the following equivalent properties holds: (1) the corresponding tautological sheaf $\mathcal{O}_{\mathbb{P}(\mathcal{E})}(1)$ is nef; (2) all its quotients have non-negative degree.
@@
The Hodge bundle of a fibration $f\colon S\to B$ can be decomposed in two ways as a direct summand of vector sub-bundles as follows:
@@
    where $\mathcal{A}$ is ample and $\mathcal{U}$ is unitary flat.
@@
We see from the above results that $q_{f}$ is the rank of the biggest trivial subbundle of $\mathcal{E}$. Moreover, we see that $\chi_{f}=\deg\mathcal{G}=\deg\mathcal{A}$. We can also derive directly from the first decomposition that $q_{f}\leq g$ and that if $q_{f}=g$ then $\chi_{f}=0$, and so $f$ is locally trivial (but Beauville’s result (4) of Theorem 2.2 is stronger).
@@
and then it holds that $q_{f}\leq u_{f}$. Observe that $\mathcal{U}$ is the biggest degree 00 sub-bundle of $f_{*}\omega_{f}$, so $u_{f}=g$ if and only if $\chi_{f}=0$ (equivalently $f$ is locally trivial): compare with (2) of Theorem 2.2.
One of the most important areas of study in this field is the so-called geography of fibred surfaces: the study of the inequalities holding between the invariants. We will list three geographical inequalities. For very nice expositions of these problems see [30] and [6].
2.1 Inequalities between the invariants: bounds on the slope $K_{f}^{2}/\chi_{f}$
Definition 2.10.
Let $f\colon S\to B$ be a locally non-trivial fibred surface. The slope of $f$ is the ratio $K_{f}^{2}/\chi_{f}$.
Usually in the literature the slope inequalities are the lower bounds on the slope.
We start by recalling two upper bounds. First of all, from Noether’s relation (2.1) and ((3)) of Theorem 2.2 we have:
@@ 2.1 Inequalities between the invariants: bounds on the slope $K_{f}^{2}/\chi_{f}$
Moreover, suppose that $f$ is relatively minimal and locally non-trivial. Then the equality in (2.12) holds if and only if $f$ is a Kodaira fibration.
@@
Given any relatively minimal locally non-trivial fibred surface $f\colon S\to B$ over a base of genus $b\geq 2$, the surface $S$ is minimal and of general type. So, the Bogomolov-Miyaoka-Yau inequality holds: $K_{S}^{2}\leq 9\chi(\mathcal{O}_{S})$. This implies that
$K_{f}^{2}\leq 9\chi_{f}+(g-1)(b-1).$
@@
The ratio 8 is particularly significant for the geometry of the surface. Indeed, consider the topological index $\tau(S)$ of the surface, i.e., the index of the intersection form on $H^{2}(S,\mathbb{C})$ (the difference between the number of its positive and negative eigenvalues). We have that $\tau(S)=\frac{1}{3}(K_{S}^{2}-2c_{2}(S))$. Therefore, using the equalities $K_{f}^{2}=K_{S}^{2}-8(g-1)(b-1)$ and $\chi_{f}:=\chi(\mathcal{O}_{S})-(g-1)(b-1)$ and the Noether equality $12\chi(\mathcal{O}_{s})=K_{S}^{2}+c_{2}(S)$, we have:
$K_{f}^{2}-8\chi_{f}=K_{S}^{2}-8\chi(\mathcal{O}_{S})=\frac{K_{S}^{2}-2c_{2}(S)}{3}=\tau(S).$
Concerning the lower bounds on the slope, the first and most famous result is the so-called classical slope inequality:
@@
[Xiao [41], Cornalba-Harris [13, 38], see also [2, Thm. 4.5], Konno [22]]
@@
If $f$ is locally non-trivial and the equality in (2.16) is reached, then $F$ is hyperelliptic and $q_{f}=0$.
@@
Even from the classical results of Theorem 2.15, it is natural to ask for lower bounds on the slope increasing with $c_{f}$ and with $q_{f}$. Many authors have contributed to this topic; here we just recall [23], [4], [5], [25].
2.2 Inequalities between the invariants: Xiao’s type of bounds
As mentioned before, we have that $q_{f}\leq g$, and the equality holds if and only if $f$ is trivial. Therefore, it is natural to ask: if $f$ is not trivial, what is the inequality between $g$ and $q_{f}$? From now on we assume $f$ is non-trivial. Xiao initiated the study in [42] by proving that if $b=0$